# A Framework for Quantum Simulations of Energy-Loss and Hadronization in Non-Abelian Gauge Theories: SU(2) Lattice Gauge Theory in 1+1D

### v2

@@ IV.3 Implementing Domain Decomposition
$\hat{O}_{M2}\equiv(\hat{X}\hat{Z}^{7}\hat{Y}-\hat{Y}\hat{Z}^{7}\hat{X})\hat{I}+\hat{I}(\hat{X}\hat{Z}^{7}\hat{Y}-\hat{Y}\hat{Z}^{7}\hat{X})\ ,\ \hat{O}_{B2}\equiv-4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{Z}^{6}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.})\ \ ,$ (37)
@@
In $\hat{O}_{Md}$, $d$ indicates that there are $d^{2}+d+1$ number of $\hat{Z}$s in between the $\hat{X}$ and $\hat{Y}$; in $\hat{O}_{Bd}$, the number of $\hat{Z}$s is $3d$ for even $d$ and 3d13d-1 for odd $d$. As preparing the ground state in the $n_{Q}=0$ sector is not the focus of this work, we do not provide details or discussions (see App. C), and focus on the $n_{Q}=1$ sector.
@@
As the $L=3$ target wavefunction differs from the tensor product of the $L=2$ and $L=1$ wavefunctions, with modifications by the boundary that penetrate into both sectors determined by the confinement scale, the fit angles in both sectors are re-optimized layer-by-layer. [^8] Such re-optimizations in the bulk of the sectors will become decreasingly significant with increasing lattice subvolumes, as the fit parameters asymptotically converge to fixed-point values in a confining theory [Klco:2019yrb, Klco:2020aud, Farrell:2023fgd, Farrell:2024fit]. This is followed by optimizing boundary operators that entangle the two sectors. Through optimization, we have identified a variational wavefunction that is prepared by a unitary operator defined in terms of 10 angles,
@@ IV.3.1 Color Complexity
Because of the key role of color entanglement in non-Abelian systems, we further explore its structure among the heavy quark and the light quarks, along with the quantum complexity measured by stabilizer Rényi entropies (SRE) [Leone:2021rzd, Haug:2022vpg, Haug:2023hcs, Leone:2024lfr, Bittel:2025yhq]. Of particular interest is the behavior of entanglement as a function of distance from the heavy quark to identify long-range quantum correlations, for which we chose to compute the mutual information $I^{(0,x)}_{f}$, which contains both classical and quantum correlations, and the 4-tangle [Wong:2000cmz], $\tau_{4f}^{(0,x)}$, for $f=q,\overline{q}$ (light quarks and anti-quarks) separately.
@@
A third quantity that we have examined is the quantum complexity, or quantum magic, measured by the SRE, ${\cal M}_{2}$ [Leone:2021rzd, Haug:2022vpg, Haug:2023hcs, Leone:2024lfr, Bittel:2025yhq]. This provides a measure of the non-stabilizerness of a quantum state [gottesman1998a, Aaronson_2004, Bravyi_2016], and how difficult it is for a classical computer to prepare it. A non-zero value of an SRE indicates that a state lies outside of those preparable by the Clifford gate set, and requires the inclusion of T-gates in quantum circuits. It has recently been shown that the evolution of magic during Quantum Approximation Optimization Algorithm (QAOA) presents a complexity barrier, as the magic increases then decreases as the target state is approached with increasing numbers of unitary layers [Capecci:2025ull]. To search for a complexity barrier in creating the ground states of the systems considered in this work, we compute the magic at each layer of operators creating the $L=3$, $n_{Q}=1$ ground state using DDec. For a system of $N$ qubits, the SRE, ${\cal M}_{\xi}$, is defined as [Leone:2021rzd]
@@
where $\hat{P}$ is a string of Pauli operators, and the sum extends over all such Pauli strings for $N$ qubits, denoted by $\hat{P}_{N}$. [^9] For $\xi=2$, and writing the wavefunction in terms of eigenstates of the $\hat{Z}$ operator with eigenvalues $\pm 1$, $|\psi\rangle=\sum_{\bm{\sigma}}c_{\bm{\sigma}}|{\bm{\sigma}}\rangle$, the expression in Eq. (42) reduces to a 4-replica summation [Tarabunga:2023xmv], $e^{-{\cal M}_{2}}=\sum_{{\bm{\sigma}}^{(1)},{\bm{\sigma}}^{(2)},{\bm{\sigma}}^{(3)},{\bm{\sigma}}^{(4)}}c_{{\bm{\sigma}}^{(1)}}\ c_{{\bm{\sigma}}^{(2)}}\ c_{{\bm{\sigma}}^{(3)}}\ c_{{\bm{\sigma}}^{(1)}{\bm{\sigma}}^{(2)}{\bm{\sigma}}^{(3)}}\ c^{*}_{{\bm{\sigma}}^{(1)}{\bm{\sigma}}^{(2)}{\bm{\sigma}}^{(4)}}\ c^{*}_{{\bm{\sigma}}^{(1)}{\bm{\sigma}}^{(3)}{\bm{\sigma}}^{(4)}}\ c^{*}_{{\bm{\sigma}}^{(2)}{\bm{\sigma}}^{(3)}{\bm{\sigma}}^{(4)}}\ c^{*}_{{\bm{\sigma}}^{(4)}}\ ,$ (43) where products of ${\bm{\sigma}}^{(s)}$ denote direct element by element multiplies. This expression provides a parametric acceleration of evaluations in sparse wavefunctions, and is what we implemented in this work (via statistical sampling). The summation in Eq. (42) scales as $\sim 4^{n_{Q}}$ for $n_{Q}$ qubits, independent of the number of non-zero amplitudes contributing to $|\psi\rangle$. The properties of the Pauli strings can be used to greatly reduce the number of evaluations. In contrast, the summation in Eq. (43) scales as $\sim{\cal N}_{\cal A}^{4}$, where ${\cal N}_{\cal A}$ is the number of non-zero amplitudes. The values of ${\cal M}_{2}$ for the $L=3$ system as a function of the variational layer are shown in Fig. 7 and given in Table 6.
@@ IV.4.1 Through the Lattice Vacuum
The exact (classically-determined) ground state of the $L=3$ system is prepared with a heavy-quark positioned at $x=0$. At $t=0$, the heavy-quark is instantaneously moved from $x=0$ to $x=1$ via the FSWAP operations defined in Eq. (22). The system is then evolved under the exact Hamiltonian in discrete time steps, but without Trotterization with respect to the individual terms contributing to the Hamiltonian. Using the same Hamiltonian parameters, $g=1.0$ and $m_{q}=0.1$, and a time step of $\delta t=0.2$, the different contributions to $\hat{H}$, as well as their sum, are computed as a function of time and are displayed in the left panel of Fig. 8. The change of heavy-quark position at $t=0$ by one spatial site injects energy into the system, by increasing the energy in the gauge field (at that instant). [^10] In the infinite-volume and continuum limits, the magnitude of the discontinuity vanishes as $g\rightarrow 0$ ($a\rightarrow 0$, and the correlation length become infinitely large (in lattice units)), and with the boundaries infinitely far away from the heavy-quark position(s), the net energy injected vanishes by Lorentz invariance. In contrast, as considered in our previous work [Farrell:2024mgu], in the presence of matter, the net injected energy will not vanish, distributed into internal excitations and fragmented hadronic particles. The extent to which the injected energy remains near the heavy quark in the form of stable hadronic excitation, or is radiated into propagating color-singlet states requires simulations of larger systems, and both effects are expected to vanish in the continuum and infinite-volume limits in the absence of a background medium.
@@ IV.4.2 In-Medium: Energy Loss and $\frac{dE}{dx}$
The $L=3$ system, despite its limitations, permits a first exploration of the energy loss in the presence of matter. To address this, and to estimate the instantaneous energy loss, $dE/dx$, in these systems with matter present, we perform two classical simulations, following the method presented in Ref. [Farrell:2024mgu]. Two systems are prepared: the ground state of the system with one heavy quark at $x=0$, and that with two heavy quarks, one at $x=0$ and the other at $x=2$. In each, the heavy quark is moved from $x=0$ to $x=1$ at $t=0$ and then to $x=2$ and $t=5$, the latter time being arbitrarily selected. The total energy of each system is determined in the three time intervals, ($t<0$, $0\leq t<5$ and $t\geq 5$) from which the difference between them can be translated into an estimator for $dE/dx$. The contributions to the energy from components of the Hamiltonian, and the total energy as a function of time are shown in Fig. 10.
@@ IV.5 Large-$L$ Tensor-Network Calculations
In order to go beyond lattice sizes limited by state vector simulations, we use tensor network techniques to study the static and dynamical properties of systems with heavy quarks. We represent the states as matrix product states (MPS), and use the density matrix renormalization group (DMRG) algorithm to prepare them in the correct sector, and the time dependent variational principle (TDVP) algorithm to perform time evolution. This is implemented via the ITensor [fishman2022itensor] and ITensorMPOContruction [Corbett:2025flm] packages. While for ground-state preparation and static property studies a maximum bond dimension of 200 was sufficient (the variance $\langle\psi|\hat{H}^{2}|\psi\rangle-\langle\psi|\hat{H}|\psi\rangle^{2}$ was below $10^{-5}$), for time evolution it had to be increased to 600. The real-time dynamics was found to be considerably harder to simulate than the static properties.
@@
By positioning the heavy quark at the center of the lattice, the energy gap to the vacuum state can be extracted with exponential convergence in lattice size, as shown in Fig. 11(b). Despite working with OBCs, where $1/L$ volume effects are present for typical quantities, this leading contribution cancels in the difference, as the heavy quark (and the screening light quarks) are sufficiently far from the boundaries. This exponential decay of correlations is also seen in the mutual information and 4-tangles as a function of distance from the heavy quark. The 4-tangles $\tau^{(x_{Q},x_{Q}+x)}_{4f}$ ($f=q,\bar{q}$), given in Eq. (41), are shown in Fig. 11(c) for a selection of (large) lattice volumes as a function of (spatial) distance to the heavy quark. As lattice volumes increase, the correlation lengths converge. Figure 11(d) shows the ${\cal M}_{2}/L$ SRE densities for both vacuum and heavy-quark states, computed via the perfect Pauli sampling method [Lami:2023naw, Haug:2023hcs] using 2000 samples. These densities are seen to converge with increasing volume, consistent with the heavy hadron having only limited spatial support on the lattice. For the small $L$ values, as explained in Ref. [Hoshino:2025ine], systems with OBCs should display a logarithmic dependence of $\mathcal{M}_{2}$ with $L$, and that is case, as seen in Fig. 11(d).
@@ V.1 Algorithms, Techniques and Quantum Circuits
There are two types of operators involved in the state preparation and time evolution: those with two $\hat{\sigma}^{\pm}$s, and those with four $\hat{\sigma}^{\pm}$s, with $\hat{Z}$s inserted in between to accommodate the fermion-exchange anti-symmetry. In the two $\hat{\sigma}^{\pm}$s case, the form $(\hat{\sigma}^{-}\hat{Z}^{n}\hat{\sigma}^{+}-{\rm h.c.})$ is used in the mesonic operators in state preparation (to furnish real wavefunctions), while $(\hat{\sigma}^{-}\hat{Z}^{n}\hat{\sigma}^{+}+{\rm h.c.})$ is used in time evolution from the kinetic term. Operators with four $\hat{\sigma}^{\pm}$s can be implemented using GHZ transformations, as presented in Refs. [Stetina:2020abi, Farrell:2022wyt]. Baryonic operators of the form $\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.}$ are used in state preparation, and those of the form $\hat{\sigma}^{+}\hat{\sigma}^{-}\hat{\sigma}^{-}\hat{\sigma}^{+}+{\rm h.c.}$ are used in time evolution from the gauge Hamiltonian, implemented by GHZ transformation circuits. The GHZ transformation used to implement the gauge Hamiltonian evolution is also used in measuring the energies.
@@ V.1.1 R-box Constructions
Building upon the work in Refs. [Algaba:2023enr, Farrell:2023fgd, Farrell:2024fit], the circuits for operators with two $\hat{\sigma}^{\pm}$s can be implemented for nearest-neighbor connectivity with two-qubit blocks, referred to as R-boxes, defined as
@@ V.1.2 State Preparation
$\hat{O}_{B0}=4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.})\ ,\ \hat{O}_{B1}=-4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{Z}^{2}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.})\ \ .$ (51)
The circuit for ${\exp}(-i\theta\hat{O}_{B0})$ is shown in Fig. 14, and ${\exp}(-i\theta\hat{O}_{B1})$ can be implemented with additional CNOTs connecting the circuit with the two qubits in the middle, as shown in Fig. 24 and discussed in App. E.
Figure 14: A circuit implementing $e^{-i\theta\hat{O}_{B0}}$, given in Eq. (51), where $G$ and $G^{\dagger}$ implement a GHZ transformation. Additional details about constructing this circuit, and that for $e^{-i\theta\hat{O}_{B1}}$, can be found in App. E.
The CNOT depth of $\hat{U}_{B0}$ is 14, while that of $\hat{U}_{B1}$ is 18.
@@
The CNOT depth of the state-preparation circuit is depth 2 for the SC vacuum (with 3 CNOT gates) for a lattice of any size, is depth 64 for the DDec of the $L=2$ sublattice and depth 20 for the $L=1$ sublattice (which can be performed in parallel). The boundary operators require an additional depth of 38, which can be partially performed in parallel. The total CNOT depth for state preparation of the $L=3$ system with $n_{Q}=1$ is depth 74.

### v3

@@ IV.3 Implementing Domain Decomposition
$\hat{O}_{M2}\equiv(\hat{X}\hat{Z}^{7}\hat{Y}-\hat{Y}\hat{Z}^{7}\hat{X})\hat{I}+\hat{I}(\hat{X}\hat{Z}^{7}\hat{Y}-\hat{Y}\hat{Z}^{7}\hat{X})\ ,\ \hat{O}_{B2}\equiv-4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{I}^{6}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.})\ \ ,$ (37)
@@
In $\hat{O}_{Md}$, $d$ indicates that there are $d^{2}+d+1$ number of $\hat{Z}$s in between the $\hat{X}$ and $\hat{Y}$; in $\hat{O}_{Bd}$, the number of $\hat{I}$s is $3d$ for even $d$ and 3d13d-1 for odd $d$. As preparing the ground state in the $n_{Q}=0$ sector is not the focus of this work, we do not provide details or discussions (see App. C), and focus on the $n_{Q}=1$ sector.
@@
As the $L=3$ target wavefunction differs from the tensor product of the $L=2$ and $L=1$ wavefunctions, with modifications by the boundary that penetrate into both sectors determined by the confinement scale, the fit angles in both sectors are re-optimized layer-by-layer. [^8] Such re-optimizations in the bulk of the sectors will become decreasingly significant with increasing lattice subvolumes, as the fit parameters asymptotically converge to fixed-point values in a confining theory [136, 137, 48, 53]. This is followed by optimizing boundary operators that entangle the two sectors. Through optimization, we have identified a variational wavefunction that is prepared by a unitary operator defined in terms of 10 angles,
@@ IV.3.1 Color Complexity
Because of the key role of color entanglement in non-Abelian systems, we further explore its structure among the heavy quark and the light quarks, along with the quantum complexity measured by stabilizer Rényi entropies (SRE) [138, 139, 140, 141, 142]. Of particular interest is the behavior of entanglement as a function of distance from the heavy quark to identify long-range quantum correlations, for which we chose to compute the mutual information $I^{(0,x)}_{f}$, which contains both classical and quantum correlations, and the 4-tangle [143], $\tau_{4f}^{(0,x)}$, for $f=q,\overline{q}$ (light quarks and anti-quarks) separately.
@@
A third quantity that we have examined is the quantum complexity, or quantum magic, measured by the SRE, ${\cal M}_{2}$ [138, 139, 140, 141, 142]. This provides a measure of the non-stabilizerness of a quantum state [144, 145, 146], and how difficult it is for a classical computer to prepare it. A non-zero value of an SRE indicates that a state lies outside of those preparable by the Clifford gate set, and requires the inclusion of T-gates in quantum circuits. It has recently been shown that the evolution of magic during Quantum Approximation Optimization Algorithm (QAOA) presents a complexity barrier, as the magic increases then decreases as the target state is approached with increasing numbers of unitary layers [147]. To search for a complexity barrier in creating the ground states of the systems considered in this work, we compute the magic at each layer of operators creating the $L=3$, $n_{Q}=1$ ground state using DDec. For a system of $N$ qubits, the SRE, ${\cal M}_{\xi}$, is defined as [138]
@@
where $\hat{P}$ is a string of Pauli operators, and the sum extends over all such Pauli strings for $N$ qubits, denoted by $\hat{P}_{N}$. [^9] For $\xi=2$, and writing the wavefunction in terms of eigenstates of the $\hat{Z}$ operator with eigenvalues $\pm 1$, $|\psi\rangle=\sum_{\bm{\sigma}}c_{\bm{\sigma}}|{\bm{\sigma}}\rangle$, the expression in Eq. (42) reduces to a 4-replica summation [148], $e^{-{\cal M}_{2}}=\sum_{{\bm{\sigma}}^{(1)},{\bm{\sigma}}^{(2)},{\bm{\sigma}}^{(3)},{\bm{\sigma}}^{(4)}}c_{{\bm{\sigma}}^{(1)}}\ c_{{\bm{\sigma}}^{(2)}}\ c_{{\bm{\sigma}}^{(3)}}\ c_{{\bm{\sigma}}^{(1)}{\bm{\sigma}}^{(2)}{\bm{\sigma}}^{(3)}}\ c^{*}_{{\bm{\sigma}}^{(1)}{\bm{\sigma}}^{(2)}{\bm{\sigma}}^{(4)}}\ c^{*}_{{\bm{\sigma}}^{(1)}{\bm{\sigma}}^{(3)}{\bm{\sigma}}^{(4)}}\ c^{*}_{{\bm{\sigma}}^{(2)}{\bm{\sigma}}^{(3)}{\bm{\sigma}}^{(4)}}\ c^{*}_{{\bm{\sigma}}^{(4)}}\ ,$ (43) where products of ${\bm{\sigma}}^{(s)}$ denote direct element by element multiplies. This expression provides a parametric acceleration of evaluations in sparse wavefunctions, and is what we implemented in this work (via statistical sampling). The summation in Eq. (42) scales as $\sim 4^{n_{Q}}$ for $n_{Q}$ qubits, independent of the number of non-zero amplitudes contributing to $|\psi\rangle$. The properties of the Pauli strings can be used to greatly reduce the number of evaluations. In contrast, the summation in Eq. (43) scales as $\sim{\cal N}_{\cal A}^{4}$, where ${\cal N}_{\cal A}$ is the number of non-zero amplitudes. The values of ${\cal M}_{2}$ for the $L=3$ system as a function of the variational layer are shown in Fig. 7 and given in Table 6.
@@ IV.4.1 Through the Lattice Vacuum
The exact (classically-determined) ground state of the $L=3$ system is prepared with a heavy-quark positioned at $x=0$. At $t=0$, the heavy-quark is instantaneously moved from $x=0$ to $x=1$ via the FSWAP operations defined in Eq. (22). The system is then evolved under the exact Hamiltonian in discrete time steps, but without Trotterization with respect to the individual terms contributing to the Hamiltonian. Using the same Hamiltonian parameters, $g=1.0$ and $m_{q}=0.1$, and a time step of $\delta t=0.2$, the different contributions to $\hat{H}$, as well as their sum, are computed as a function of time and are displayed in the left panel of Fig. 8. The change of heavy-quark position at $t=0$ by one spatial site injects energy into the system, by increasing the energy in the gauge field (at that instant). [^10] In the infinite-volume and continuum limits, the magnitude of the discontinuity vanishes as $g\rightarrow 0$ ($a\rightarrow 0$, and the correlation length become infinitely large (in lattice units)), and with the boundaries infinitely far away from the heavy-quark position(s), the net energy injected vanishes by Lorentz invariance. In contrast, as considered in our previous work [118], in the presence of matter, the net injected energy will not vanish, distributed into internal excitations and fragmented hadronic particles. The extent to which the injected energy remains near the heavy quark in the form of stable hadronic excitation, or is radiated into propagating color-singlet states requires simulations of larger systems, and both effects are expected to vanish in the continuum and infinite-volume limits in the absence of a background medium.
@@ IV.4.2 In-Medium: Energy Loss and $\frac{dE}{dx}$
The $L=3$ system, despite its limitations, permits a first exploration of the energy loss in the presence of matter. To address this, and to estimate the instantaneous energy loss, $dE/dx$, in these systems with matter present, we perform two classical simulations, following the method presented in Ref. [118]. Two systems are prepared: the ground state of the system with one heavy quark at $x=0$, and that with two heavy quarks, one at $x=0$ and the other at $x=2$. In each, the heavy quark is moved from $x=0$ to $x=1$ at $t=0$ and then to $x=2$ and $t=5$, the latter time being arbitrarily selected. The total energy of each system is determined in the three time intervals, ($t<0$, $0\leq t<5$ and $t\geq 5$) from which the difference between them can be translated into an estimator for $dE/dx$. The contributions to the energy from components of the Hamiltonian, and the total energy as a function of time are shown in Fig. 10.
@@ IV.5 Large-$L$ Tensor-Network Calculations
In order to go beyond lattice sizes limited by state vector simulations, we use tensor network techniques to study the static and dynamical properties of systems with heavy quarks. We represent the states as matrix product states (MPS), and use the density matrix renormalization group (DMRG) algorithm to prepare them in the correct sector, and the time dependent variational principle (TDVP) algorithm to perform time evolution. This is implemented via the ITensor [149] and ITensorMPOContruction [150] packages. While for ground-state preparation and static property studies a maximum bond dimension of 200 was sufficient (the variance $\langle\psi|\hat{H}^{2}|\psi\rangle-\langle\psi|\hat{H}|\psi\rangle^{2}$ was below $10^{-5}$), for time evolution it had to be increased to 600. The real-time dynamics was found to be considerably harder to simulate than the static properties.
@@
By positioning the heavy quark at the center of the lattice, the energy gap to the vacuum state can be extracted with exponential convergence in lattice size, as shown in Fig. 11(b). Despite working with OBCs, where $1/L$ volume effects are present for typical quantities, this leading contribution cancels in the difference, as the heavy quark (and the screening light quarks) are sufficiently far from the boundaries. This exponential decay of correlations is also seen in the mutual information and 4-tangles as a function of distance from the heavy quark. The 4-tangles $\tau^{(x_{Q},x_{Q}+x)}_{4f}$ ($f=q,\bar{q}$), given in Eq. (41), are shown in Fig. 11(c) for a selection of (large) lattice volumes as a function of (spatial) distance to the heavy quark. As lattice volumes increase, the correlation lengths converge. Figure 11(d) shows the ${\cal M}_{2}/L$ SRE densities for both vacuum and heavy-quark states, computed via the perfect Pauli sampling method [151, 140] using 2000 samples. These densities are seen to converge with increasing volume, consistent with the heavy hadron having only limited spatial support on the lattice. For the small $L$ values, as explained in Ref. [152], systems with OBCs should display a logarithmic dependence of $\mathcal{M}_{2}$ with $L$, and that is case, as seen in Fig. 11(d).
@@ V.1 Algorithms, Techniques and Quantum Circuits
There are two types of operators involved in the state preparation and time evolution: those with two $\hat{\sigma}^{\pm}$s, with $\hat{Z}$s inserted in between to accommodate the fermion-exchange anti-symmetry, and those with four $\hat{\sigma}^{\pm}$s. In the two $\hat{\sigma}^{\pm}$s case, the form $(\hat{\sigma}^{-}\hat{Z}^{n}\hat{\sigma}^{+}-{\rm h.c.})$ is used in the meson operators in state preparation (to furnish real wavefunctions), while $(\hat{\sigma}^{-}\hat{Z}^{n}\hat{\sigma}^{+}+{\rm h.c.})$ is used in time evolution from the kinetic term. Operators with four $\hat{\sigma}^{\pm}$s can be implemented using GHZ transformations, as presented in Refs. [153, 40]. Baryon operators of the form $\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.}$ are used in state preparation, and those of the form $\hat{\sigma}^{+}\hat{\sigma}^{-}\hat{\sigma}^{-}\hat{\sigma}^{+}+{\rm h.c.}$ are used in time evolution from the gauge Hamiltonian, implemented by GHZ transformation circuits. The GHZ transformation used to implement the gauge Hamiltonian evolution is also used in measuring the energies.
@@ V.1.1 R-box Constructions
Building upon the work in Refs. [133, 48, 53], the circuits for operators with two $\hat{\sigma}^{\pm}$s can be implemented for nearest-neighbor connectivity with two-qubit blocks, referred to as R-boxes, defined as
@@ V.1.2 State Preparation
$\hat{O}_{B0}=4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.})\ ,\ \hat{O}_{B1}=-4i(\hat{\sigma}^{+}\hat{\sigma}^{+}\hat{I}^{2}\hat{\sigma}^{-}\hat{\sigma}^{-}-{\rm h.c.})\ \ .$ (51)
The circuit for ${\exp}(-i\theta\hat{O}_{B0})$ is shown in Fig. 14, and ${\exp}(-i\theta\hat{O}_{B1})$ can be implemented with two additional qubits in the middle. Additional discussions can be found in App. E.
Figure 14: A circuit implementing $e^{-i\theta\hat{O}_{B0}}$, given in Eq. (51), where $G$ and $G^{\dagger}$ implement a GHZ transformation. Additional details about constructing this circuit can be found in App. E.
The CNOT depth of $\hat{U}_{Bd}$ on an all-to-all connected quantum device (all-to-all CNOT depth) is 14.
@@
The all-to-all CNOT depth of the state-preparation circuit is depth 2 for the SC vacuum (with 3 CNOT gates) for a lattice of any size, is depth 50 for the DDec of the $L=2$ sublattice with cancellations and depth 20 for the $L=1$ sublattice (which can be performed in parallel). The boundary operators require an additional depth of 38, which can be partially performed in parallel. The total all-to-all CNOT depth for state preparation of the $L=3$ system with $n_{Q}=1$ is depth 72.